Suppose two angles are supplements. Fine the measures of both angles if one of the angles is 69 degrees smaller than twice the other angle.
I understamd what to do but if the first problem equals to 180 what does the second equation equal to?

Answer by rapaljer(4671)

Let y="other angle"
x= "one of the angles"

x+y=180
x=2y-69

Solve by substitution:
x+y=180
(2y-69) + y = 180
3y - 69 = 180
3y = 249
y = 83
x=97

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida

www.mathinlivingcolor.com.